OpenStudy (anonymous):

Before taking a typing course, Frank could type 27 words per minute by the end of the course he was able to type 77 words per minute. Find the percent increase. A 185.2% B 74% C 35.1% D 64.9%

OpenStudy (anonymous):

% Change is ALWAYS (end point - starting point) / (starting point) end point is 77 starting point is 27 % change = (77-27) / 27 = 50 / 27 = 1.85, or 1.85 X 100 = 185% Since he increased his typing speed, it's correct to call the % change a % increase A brief word about our calculation: For 50 / 27 we got 1.85, and that's the correct answer. Note that 1.85 is really 1.85 / 1 Since they want % change, we need to remember that the % sign stands for "per hundred", and we only have 1.85 / 1 , or per one. Simply multiply top and bottom of 1.85 / 1 by 100 to get 185/100, or 185%